OpenStudy (anonymous):

2a/3+1/5=7/6 (meant to be fractions..)

OpenStudy (anonymous):

2a/3+6/30=35/30 2a/3=29/30 20a/30=29/30 20a=29 a=1.45

OpenStudy (agreene):

(2 a)/3+1/5 = 7/6 Subtract 1/5 from both sides: (2 a)/3 = 29/30 Divide both sides by 2/3: a = 29/20